OpenStudy (anonymous):

Complete the ordered pair of (0,_) with y=-3x+7

OpenStudy (dumbcow):

For every ordered pair (x-value, y-value) -> (0,y) y = -3(x-value)+7 plug in 0 for x y=-3*0 +7=0+7=7